Shutdown pquest SNF 2004-11-22 23:24:09: chuck oring not seated

eenriquez at snf.stanford.edu eenriquez at snf.stanford.edu
Tue Nov 23 15:01:13 PST 2004


Vented chamber and installed a new silicone o-ring.  Zeroed the backside He flowmeter.  Also installed additional aluminum foil around the backside He flometer and manometer to shield it from RF noise.  Now He flow without a wafer is ~ 8 mT, with wafer = ~2.2 mT.  Ran several recipes with different power settings  with no problems.




More information about the pquest-pcs mailing list